OpenStudy (anonymous):

write an equation to represent the statement "four less then the product of three and x is twenty-three "

OpenStudy (anonymous):

3x-4=23

OpenStudy (kainui):

break it up into parts. Every time they say a word meaning something, write that symbol down. So when they say "four less than ..." that means whatever that stuff is minus four. So write -4 on your paper and keep adding it bit by bit until you have the expression.

OpenStudy (anonymous):

To make this, you can do 3x - 4 = 23, or -4 + 3x = 23. If you have to solve it for x, you can add 4 to each side, making 3x = 27, then divide each side by 3, making it x = 9.

OpenStudy (anonymous):

so whats the answer?

OpenStudy (anonymous):

answer is 3x-4=23

OpenStudy (anonymous):

no solving or anythings, its simply 3x-4=23

OpenStudy (anonymous):

you break it up into parts 4 less than means you will take four from something that something in this case is the product of 3 and x and a product is multiplying right so its product of three and x= (3x) four less than (- 4) your subtracting four 3x-4 and the final part is its telling you that equation is 23 3x-4=23

OpenStudy (anonymous):

the statement is 3x-4=23